OpenStudy (anonymous):

Rewrite the following in simplified radical form. √27w^10 Assume that all variables represent positive real numbers.

OpenStudy (anonymous):

IS THIS sqrt(27w^2)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait no, its sqrt(27w^10)

OpenStudy (anonymous):

OK http://www.wolframalpha.com/input/?i=%E2%88%9A27w^10

OpenStudy (anonymous):

nice try iv used wolframalpha before, its not simplified radical form

OpenStudy (anonymous):

Well, factoring the part under the radical we have: \[\sqrt{27w^{10}} = \sqrt{(3^2\cdot 3)(w^2\cdot w^2 \cdot w^2\cdot w^2\cdot w^2)}\] And pulling out each squared factor we get: \[3\sqrt{3}|w^5|\]

OpenStudy (anonymous):

You can also do this the quick way by recalling that \[\sqrt[b]{a^k} = (a)^{k\over b}\]